# Simulating Policy Impacts: Developing a Generative Scenario Writing Method to Evaluate the Perceived Effects of Regulation

### v1

@@ Abstract
The rapid advancement of AI technologies yields numerous future impacts on individuals and society. Policy-makers are therefore tasked to react quickly and establish policies that mitigate those impacts. However, anticipating the effectiveness of policies is a difficult task, as some impacts might only be observable in the future and respective policies might not be applicable to the future development of AI. In this work we develop a method for using large language models (LLMs) to evaluate the efficacy of a given piece of policy at mitigating specified negative impacts. We do so by using GPT-4 to generate scenarios both pre- and post-introduction of policy and translating these vivid stories into metrics based on human perceptions of impacts. We leverage an already established taxonomy of impacts of generative AI in the media environment to generate a set of scenario pairs both mitigated and non-mitigated by the transparency legislation of Article 50 of the EU AI Act (Commission 2024). We then run a user study ($n=234$) to evaluate these scenarios across four risk-assessment dimensions: severity, plausibility, magnitude, and specificity to vulnerable populations. We find that this transparency legislation is perceived to be effective at mitigating harms in areas such as labor and well-being, but largely ineffective in areas such as social cohesion and security. Through this case study on generative AI harms we demonstrate the efficacy of our method as a tool to iterate on the effectiveness of policy on mitigating various negative impacts. We expect this method to be useful to researchers or other stakeholders who want to brainstorm the potential utility of different pieces of policy or other mitigation strategies.
@@ 1 Introduction
In addition to creating a suite of positive benefits from facilitating ease of culturally situated natural language creation and translation (Yao et al. 2023), to assisting with medical imaging (Chen et al. 2022), and empowering new and accessible forms of education (Alasadi and Baiz 2023), generative AI has given reason for concern in many aspects of society. Any implementation of a system using generative AI, or any algorithmic tool for that matter, can have unintended downstream effects—both positive and negative. It is essential to have a comprehensive understanding of these potential impacts when considering how to make decisions regarding the usage of these tools.
@@
Some LLMs, such as GPT-4, have been trained on more data than any human could consume in a millennium (Achiam et al. 2023). As such, they are able to learn enough about the world in terms of knowledge and relationships to be able to (to some extent) simulate some underlying cause and effect relationships in the world (Li, Nye, and Andreas 2021; Cai, Liu, and Song 2023). We aim to utilize this underlying ‘world knowledge’ in such models to simulate the impacts of AI in society given a policy implementation such as a particular Article of the EU AI Act (Commission 2024). We do so by utilizing written scenarios as small world representations that reflect and embed social and causal constraints within their text. Specifically, we prompt GPT-4 to write a large and varied set of scenarios each in terms of a particular impact about generative AI (e.g., sensationalism in regards to media quality impacts) in the media environment to take place in the US in the next five years. We then introduce a mitigation strategy in terms of a policy implementation, and ask the model to re-write a given scenario in light of the introduced policy text. In this study we examine the impact of Article 50 of the EU AI Act about transparency obligations. To evaluate the quality of the scenarios and measure the influence of the policy on the impacts represented in the scenarios we then run a user study where we ask a set of human raters to read both the original scenario and the re-written scenario and then rate them on four dimensions that are of high relevance in impact assessment and scenario-writing literature: severity, plausibility, magnitude, and specificity to vulnerable populations.
@@ 2 Related Literature
The National Institute of Science and Technologies (NIST) introduced a structure to improve the trustworthiness of AI that follows a map, measure, and manage voluntary framework to deploy AI technologies (NIST 2023). They acknowledge the sociotechnical (Shelby et al. 2023) nature of these risks, and suggest the use of this framework to both initiate AI risk management and bolster existing systems. The idea is to (1) map out the potential harms of these AI systems; this is an active area of research identifying AI harms such as (Kieslich, Diakopoulos, and Helberger 2023; Kieslich, Helberger, and Diakopoulos 2024; Barnett 2023; Weidinger et al. 2023; Chanda and Banerjee 2022). The measuring step (2) involves methods to quantify these risks and enable assessment of whether we are mitigating them successfully. Finally, the managing step (3) involves integration of responsible policies to prevent these harms from manifesting. Our proposed method allows stakeholders to measure harms in an exploratory manner with the goal of enabling people in positions of power (e.g., legislators and government officials) to introduce policy to manage these harms.
@@
Scenario writing is one method that can be an effective approach to stimulate future thinking of laypeople (Amer, Daim, and Jetter 2013; Amos‐Binks, Dannenhauer, and Gilpin 2023; Börjeson et al. 2006; Burnam-Fink 2015; Selin 2006). The goal of scenario-writing is to create narratives about future impact of technology that vividly show how a specific technology impacts the life of story characters. Numerous plausible and potential futures emerge that can then be used as a conversation starter to develop mitigation strategies or to classify user-centric impacts (Amer, Daim, and Jetter 2013; Burnam-Fink 2015). The impact assessment literature has acknowledged the value of scenarios (Meßmer and Degeling 2023) and several studies have relied on laypeople to compose future impact scenarios (Diakopoulos and Johnson 2021; Kieslich, Diakopoulos, and Helberger 2023; Kieslich, Helberger, and Diakopoulos 2024), use scenarios as a conversation starter for debate about ethics (Das et al. 2024), or provide situations to collect judgements related to AI ethics (Awad et al. 2018).
@@ 3 Methodology
We used GPT-4111Specifically “gpt-4-turbo” available via the OpenAI API which pointed to the most recent model available as of Feb. 2024. (Achiam et al. 2023) for this study since at the time of data collection it outperformed other proprietary and open LLMs on the ChatArena benchmark which evaluates models according to pairwise human comparisons (Chiang et al. 2024) [^2] https://chat.lmsys.org/?leaderboard. We used OpenAI’s API for chat completions to generate scenarios. To generate scenarios for our study we start with a set of 50 impact types, organized into 10 higher-level impact themes, that describe various impacts of generative AI in the media and information ecosystem (Kieslich, Diakopoulos, and Helberger 2023). Future work may consider extending the methods developed here to other impact typologies and domains. We next describe how and why we arrived at our prompting approach and then elaborate on how we generated scenarios to support our study.
@@
Developing a Prompting Approach. We initially experimented with simply prompting the model with the 10 general impact themes in order to understand the diversity with which GPT-4 could uncover potential impacts when only prompted with a general impact theme. We thus prompted GPT-4 to generate scenarios for each of the ten impact themes ten times each: autonomy, education, labor, legal rights, media quality, political, security, social cohesion, trustworthiness, and well-being. In the Appendix, Figure 5 displays the number of specific impact types that were uncovered by the LLM, as well as how often (in the ten runs) various specific impact types were discussed.
@@
For all themes except trustworthiness, when using this prompting approach the model was not able to generate scenarios reflecting each of the specific impact types in the typology outlined by Kieslich, Diakopoulos, and Helberger (2023). More importantly, for each of the impact themes, the model tended to heavily skew toward discussing a couple of specific impact types and rarely touched on others. For instance, prompting only with the impact theme of well-being produced scenarios covering three out of four specific impact types from the typology, and generated scenarios discussed mental harm 90% of the time, never mentioned physical harm, and only discussed addiction and reputation two out of ten times each.
@@
Due to this skew and non-comprehensive spread of impact types present in the scenarios when prompting only with impact theme, we moved on to experiment with prompting the model explicitly for each of the 50 impact types rather than at the level of the ten impact themes. It became evident that utilizing LLMs to generate scenarios was more effective when done with a human-in-the-loop providing contextual knowledge (in this case, the full taxonomy provided by Kieslich, Diakopoulos, and Helberger (2023)) in order to guide creation of specific outputs—if we only used general impact themes we would not uncover as many impacts or the same level of detail. Thus, for the remainder of this paper, we evaluate scenarios generated by GPT-4 that were generated with inclusion of each specific impact type.
@@
We refer to the prompt for the illustrative scenario, $S$, as Prompt 1, and the prompt for the policy-mitigated scenario, $S^{\prime}$, as Prompt 2.333For each scenario pair we utilized a fresh state of the chatbot with temperature = 0.7. For Prompt 1, We first included context to define what a scenario was, name the specific impact type in the context of the impact theme, and define generative AI. We then directed the model to write a short fictional scenario set in the United States in the year 2029 about this impact, and included specific instructions for the model not to introduce reflection or meta analysis in the scenario since we reasoned that this might influence or bias the interpretation of participants in our study. For Prompt 2, we indicated to the model that a piece of legislation was enacted and that it should rewrite the scenario in light of that legislation. For this study, we provided it the exact text of the final version of Article 50 of the EU AI Act (Commission 2024), which concerns transparency obligations for certain AI systems, including for general purpose AI models.
Scenario Validation
To validate the scenarios for the study, two of the authors independently evaluated the 150 scenarios in order to assess whether the scenarios actually depicted the designated impact provided in the corresponding prompt. Evaluators rated each scenario as (1) depicting the impact, (2) somewhat depicting the impact, and (3) not depicting the impact. A scenario pair needed either both evaluators to assess it as depicting the impact or one evaluator saying it depicted and the other saying it was somewhat present to be considered valid. If zero or only one of the three scenarios were deemed valid, the specific impact type was removed from further evaluation. If two of the three scenarios passed the validation check, the third scenario was regenerated and then reassessed by both evaluators to ensure all three scenarios depicted the designated impacts.
@@
We excluded respondents that failed at least two attention checks. Our attention checks took the form of nonsensical statements for which there was a logical and common sense response (e.g., To what extent do you agree with this statement: “I’ve developed a way to photosynthesize, so I sustain myself on sunlight alone, like a plant.”? This is to check your attention.) [^4] See: Prolific’s Attention and Comprehension Check Policy. https://tinyurl.com/prolific-attention ; Our attention checks are available at: https://tinyurl.com/attention-checks. Evaluators were not allowed to take the survey multiple times. Fifteen raters were removed via attention checks (6% of total 249 recruited) resulting in a final participant pool of 234 raters total. 12% of evaluators did not consent to demographic data collection through Prolific, but the remaining sample was comprised as follows: a fairly equal dispersion of male and female identifying evaluators (51% F; 48% M; 1% prefer not to say), predominantly white (68%), followed by 11% Mixed, 8% Black, 7% Asian, and 5% other; and an average age of 38 (minimum: 19, maximum: 76).
@@
We chose to assess the content of scenarios across four dimensions to assess the impacts portrayed: severity, plausibility, magnitude, and specificity to vulnerable populations.
@@
Plausibility is an essential evaluation dimension for risk analysis (Glette-Iversen, Aven, and Flage 2022) and scenario quality (Amer, Daim, and Jetter 2013). Following the typology consolidated by Börjeson et al. (2006), we are generating explorative scenarios in this work. As such we need a measure to assess the degree to which the events in a scenario are realistic or at least consistent with known social dynamics and forces such that one could reasonably conclude that a scenario could happen based on their world knowledge (Urueña 2019).
@@
We intentionally wanted to allow evaluators’ intuition to guide them to evaluate these rather than have any sort of anchoring guide such as giving an explicit example of what would constitute ‘extremely severe’. The definitions of the items were developed from dictionary definitions in order to elicit ratings from colloquial understandings, rather than injecting expert definitions and biases into the evaluations. After a rater evaluated the first scenario, they were then presented with “a re-written version of the above scenario” (either $S$ or $S^{\prime}$, whichever they did not see first), and performed the same evaluation on this second scenario. To mitigate the potential for ordering effects we randomly counterbalance the order of presentation: three evaluators saw each scenario pair in the order $S$, $S^{\prime}$, and three people saw the scenario in the order $S^{\prime}$, $S$. We later performed a paired sample t-test to assess if there were any ordering biases and found none overall or when aggregated by impact type.
@@ 4 Analysis and Results
For each impact theme below, we will examine (a) which scenarios GPT-4 was adequately able to generate when specifically prompted, (b) the human user study raw scores to understand how these impacts were perceived, and (c) the delta between $S$ and $S^{\prime}$ to evaluate whether humans perceived the policy to be effective at mitigating the designated harm. We begin by examining overall significant trends, then proceed alphabetically by impact theme. All references to significance refer to results from paired sample t-tests, and we report significance at levels of $p<0.05,0.01$, and $0.001$. All references to a taxonomy of impacts of generative AI in the media environment uncovered by human authors refer to (Kieslich, Diakopoulos, and Helberger 2023).
@@
Now we will drill into the specific impact themes, referencing individual specific impact types (Appendix Tables 2-11) when relevant. We note significance levels when present, otherwise discussing general trends.
@@
Evaluators believed these issues affected a large portion of society ($M_{S}=4.06;SD_{S}=0.97$) with the second highest average scores for magnitude. They believe the mitigated impacts affected slightly fewer people ($M_{\Delta}=-0.17;SD_{\Delta}=1.01$), and similarity the mitigation made the impacts less specific to vulnerable populations ($M_{\Delta}=-0.44;SD_{\Delta}=0.83$).
@@
Political impacts across the board tended to be perceived as the most severe of all impact themes ($M_{S}=4.08;SD_{S}=0.76$), with fake news/misinformation having the highest severity ratings of all impact types with a score of $M_{S}=4.56$ ($SD_{S}=0.50$), and manipulation not trailing too far behind ($M_{S}=4.22;SD_{S}=0.92$). On average, transparency legislation seemed to mitigate the severity ($M_{\Delta}=-0.56$***; $SD_{\Delta}=0.86;p<0.001$) of political impacts and even more so in fake news/misinformation ($M_{\Delta}=-0.78$*; $SD_{\Delta}=0.63;p<0.05$) and manipulation ($M_{\Delta}=-0.67$*; $SD_{\Delta}=0.82;p<0.05$). One of the scenarios portraying fake news detailed an entirely fabricated political scandal concocted by an AI news generator and widely disseminated to the public that ultimately led the political candidate to drop out of the race. In the transparency mitigated scenario, an “AI-generated” tag was attached to the piece, and news consumers all paid attention to the human-verified fact checking of the inaccuracies of the purported scandal.
These scenarios detailing political impacts were seen as very plausible ($M_{S}=4.11;SD_{S}=0.97$), second only to labor scenarios. They also notably were one of the only impact themes for which the transparency legislation had a perceived significant impact on the specificity to vulnerable populations ($M_{\Delta}=-0.25$*; $SD_{\Delta}=0.68;p<0.05$), indicating that the perceived effect of this legislation may be even more beneficial to those more disparately affected by AI harms.
@@
We were able to generate adequately relevant scenarios for each of the impact types within trustworthiness: discernment of fact and fiction, information chaos, media fatigue, mistrust, and overreliance on AI. These scenarios tended to focus on the relationship between readers and journalists or readers and the news environment at large. They also tended to feature a trajectory towards a lack of trust between society and the news we consume.
@@
5 Discussion and Further Work

### v2

@@ Abstract
The rapid advancement of AI technologies yields numerous future impacts on individuals and society. Policymakers are tasked to react quickly and establish policies that mitigate those impacts. However, anticipating the effectiveness of policies is a difficult task, as some impacts might only be observable in the future and respective policies might not be applicable to the future development of AI. In this work we develop a method for using large language models (LLMs) to evaluate the efficacy of a given piece of policy at mitigating specified negative impacts. We do so by using GPT-4 to generate scenarios both pre- and post-introduction of policy and translating these vivid stories into metrics based on human perceptions of impacts. We leverage an already established taxonomy of impacts of generative AI in the media environment to generate a set of scenario pairs both mitigated and non-mitigated by the transparency policy in Article 50 of the EU AI Act. We then run a user study ($n=234$) to evaluate these scenarios across four risk-assessment dimensions: severity, plausibility, magnitude, and specificity to vulnerable populations. We find that this transparency legislation is perceived to be effective at mitigating harms in areas such as labor and well-being, but largely ineffective in areas such as social cohesion and security. Through this case study we demonstrate the efficacy of our method as a tool to iterate on the effectiveness of policy for mitigating various negative impacts. We expect this method to be useful to researchers or other stakeholders who want to brainstorm the potential utility of different pieces of policy or other mitigation strategies.
@@ 1 Introduction
In addition to creating a suite of positive benefits, from facilitating ease of culturally situated natural language creation and translation (Yao et al. 2023), to assisting with medical imaging (Chen et al. 2022), and empowering accessible forms of education (Alasadi and Baiz 2023), generative AI has given reason for concern in many aspects of society. Any implementation of a system using generative AI, or any algorithmic tool for that matter, can have unintended downstream effects. It is essential to have a comprehensive understanding of these potential impacts when considering how to make decisions regarding the usage of these tools.
@@
Some LLMs, such as GPT-4, have been trained on more data than any human could consume in a millennium (Achiam et al. 2023). As such, they are able to learn enough about the world in terms of knowledge and relationships to be able to (to some extent) simulate some underlying cause and effect relationships in the world (Li, Nye, and Andreas 2021; Cai, Liu, and Song 2023). We aim to utilize this underlying ‘world knowledge’ in such models to simulate the impacts of AI in society given a policy implementation such as a particular Article of the EU AI Act (Commission 2024). We do so by utilizing written scenarios as small world representations that reflect and embed social and causal constraints within their text. Specifically, we prompt GPT-4 to write a large and varied set of scenarios each in terms of a particular impact about generative AI (e.g., sensationalism in regards to media quality impacts) to take place in the US in the next five years. We then introduce a mitigation strategy in terms of a policy implementation, and ask the model to re-write a given scenario in light of the introduced policy text. In this study we examine the impact of Article 50 of the EU AI Act as a specific expression of a set of policy ideas related to transparency obligations. To evaluate the quality of the scenarios and measure the influence of the policy on the impacts represented in the scenarios we then run a user study asking a set of human raters to read both the original and re-written scenario and then rate them on four dimensions that are of high relevance in impact assessment (IA) and scenario-writing literature: severity, plausibility, magnitude, and specificity to vulnerable populations.
@@ 2 Related Literature
The National Institute of Science and Technologies (NIST) introduced a structure to improve the trustworthiness of AI that follows a map, measure, and manage voluntary framework to deploy AI technologies (NIST 2023). They acknowledge the sociotechnical (Shelby et al. 2023) nature of these risks, and suggest the use of this framework to both initiate AI risk management and bolster existing systems. The idea is to (1) map out the potential harms of these AI systems; this is an active area of research identifying AI harms such as (Kieslich, Diakopoulos, and Helberger 2024; Kieslich, Helberger, and Diakopoulos 2024; Barnett 2023; Weidinger et al. 2023; Chanda and Banerjee 2022). The measuring step (2) involves methods to quantify these risks and enable assessment of whether we are mitigating them successfully. Finally, the managing step (3) involves integration of responsible policies to prevent these harms from manifesting. Our proposed method allows stakeholders to measure harms in an exploratory manner with the goal of enabling people in positions of power (e.g., legislators and government officials) to introduce policy to manage these harms.
@@
Scenario writing is one method that can be an effective approach to stimulate future thinking of laypeople (Amer, Daim, and Jetter 2013; Amos‐Binks, Dannenhauer, and Gilpin 2023; Börjeson et al. 2006; Burnam-Fink 2015; Selin 2006). The goal of scenario-writing is to create narratives about future impact of technology that vividly show how a specific technology impacts the life of story characters. Numerous plausible and potential futures emerge that can then be used as a conversation starter to develop mitigation strategies or to classify user-centric impacts (Amer, Daim, and Jetter 2013; Burnam-Fink 2015). The impact assessment literature has acknowledged the value of scenarios (Meßmer and Degeling 2023) and several studies have relied on laypeople to compose future impact scenarios (Diakopoulos and Johnson 2021; Kieslich, Diakopoulos, and Helberger 2024; Kieslich, Helberger, and Diakopoulos 2024), use scenarios as a conversation starter for debate about ethics (Das et al. 2024), or provide situations to collect judgements related to AI ethics (Awad et al. 2018).
@@ 3 Methodology
We used GPT-4111“gpt-4-turbo” available via the OpenAI API in Feb. 2024. (Achiam et al. 2023) for this study since at the time of data collection it outperformed other proprietary and open LLMs on the ChatArena benchmark which evaluates models according to pairwise human comparisons (Chiang et al. 2024) [^2] https://chat.lmsys.org/?leaderboard. We used OpenAI’s API for chat completions to generate scenarios. To generate scenarios illustrating various impacts we ground the approach with a set of 50 impact types, organized into 10 higher-level impact themes that describe various impacts of generative AI in the media and information ecosystem (Kieslich, Diakopoulos, and Helberger 2024). This “Kieslich taxonomy” as we refer to it here, was itself developed by analyzing scenarios written by a diverse array of respondents to an anticipatory scenario writing task. The 10 impact themes identified include: autonomy, education, labor, legal rights, media quality, political, security, social cohesion, trustworthiness, and well-being, and within these there are 50 specific impact types which we explicitly use in our scenario generation process and refer to in our findings. By leveraging an already established taxonomy of specific impacts, we not only ground the approach and addresses plausibility and bias issues created by more speculative LLM use, but also have an established thematic grouping to guide our analysis and interpretation. Future work may consider extending the methods developed here to other impact typologies and domains.
@@
Developing a Prompting Approach. We initially experimented with simply prompting the model with the 10 general impact themes in the Kieslich taxonomy in order to understand the diversity with which GPT-4 could uncover potential impacts when only prompted with a general impact theme. We thus prompted GPT-4 to generate scenarios for each of the ten impact themes ten times each. In the Appendix, Figure 5 displays the number of specific impact types that were uncovered by the LLM, as well as how often various specific impact types were discussed.
@@
When using this prompting approach the model was not able to generate scenarios reflecting each of the specific impact types outlined in the Kieslich taxonomy except for trustworthiness. More importantly, for each of the impact themes, the model tended to heavily skew toward discussing a couple of specific impact types and rarely touched on others. For instance, prompting only with the impact theme of well-being produced scenarios covering three out of four specific impact types from the typology, and generated scenarios discussed mental harm 90% of the time, never mentioned physical harm, and only discussed addiction and reputation two out of ten times each.
@@
Due to this skew and non-comprehensive spread of impact types present in the scenarios when prompting only with impact theme, we moved on to experiment with prompting the model explicitly for each of the 50 impact types rather than at the level of the ten impact themes. It became evident that utilizing LLMs to generate scenarios was more effective when done with a human-in-the-loop providing contextual knowledge (in this case, the full Kieslich taxonomy) in order to guide creation of specific outputs—if we only used general impact themes we would not uncover as many impacts or the same level of detail. Thus, for the remainder of this paper, we evaluate scenarios generated by GPT-4 that were generated with inclusion of each specific impact type.
@@
We refer to the prompt for the illustrative scenario, $S$, as Prompt 1, and the prompt for the policy-mitigated scenario, $S^{\prime}$, as Prompt 2.333For each scenario pair we used a fresh state with temp. 0.7. For Prompt 1, We first included context to define what a scenario was, name the specific impact type in the context of the impact theme, and define generative AI. We then directed the model to write a short fictional scenario set in the United States in the year 2029 about this impact, and included specific instructions for the model not to introduce reflection or meta analysis in the scenario since we reasoned that this might influence or bias the interpretation of participants in our study. For Prompt 2, we indicated to the model that a piece of legislation was enacted and that it should rewrite the scenario in light of that legislation. For this study, we provided it the exact text of the final version of Article 50 of the EU AI Act (Commission 2024), which concerns transparency obligations for certain AI systems, including for general purpose AI models.
Scenario Validation.
To assess the extent to which GPT-4 adequately generated scenarios for each of the impact types, we had two authors independently validate each of the 150 scenarios in order to assess whether the scenarios actually depicted the designated impact provided in the corresponding prompt. Evaluators rated each scenario as (1) depicting the impact, (2) somewhat depicting the impact, and (3) not depicting the impact. A scenario pair needed either both evaluators to assess it as depicting the impact or one evaluator saying it depicted and the other saying it was somewhat present to be considered valid. If zero or only one of the three scenarios were deemed valid, the specific impact type was removed from further evaluation. If two of the three scenarios passed the validation check, the third scenario was regenerated and then reassessed by both evaluators to ensure all three scenarios depicted the designated impacts.
@@
We excluded respondents that failed at least two attention checks. Our attention checks took the form of nonsensical statements for which there was a logical and common sense response (e.g., To what extent do you agree with this statement: “I’ve developed a way to photosynthesize, so I sustain myself on sunlight alone, like a plant.”? This is to check your attention.) [^4] See: Prolific’s Attention and Comprehension Check Policy. https://tinyurl.com/prolific-attention ; Our attention checks are available at: https://tinyurl.com/attention-checks. Evaluators were not allowed to take the survey multiple times. Fifteen raters were removed via attention checks (6% of 249 recruited) resulting in a final participant pool of 234 raters. 12% of evaluators did not consent to demographic data collection through Prolific, but the remaining sample was comprised as follows: a fairly equal dispersion of male and female identifying evaluators (51% F; 48% M; 1% prefer not to say), predominantly white (68%), followed by 11% Mixed, 8% Black, 7% Asian, and 5% other; and an average age of 38 (minimum: 19, maximum: 76).
@@
We chose to assess the impacts portrayed in the scenarios across four dimensions: severity, plausibility, magnitude, and specificity to vulnerable populations.
@@
Plausibility is an essential evaluation dimension for risk analysis (Glette-Iversen, Aven, and Flage 2022) and scenario quality (Amer, Daim, and Jetter 2013). Following the typology consolidated by Börjeson et al. (2006), we are generating explorative scenarios in this work; we need a measure to assess the degree to which the events in a scenario are realistic or consistent with known social dynamics such that one could reasonably conclude that a scenario could happen based on their world knowledge (Urueña 2019).
@@
We wanted to allow evaluators’ intuition to guide them to evaluate these rather than have any sort of anchoring guide such as giving an explicit example of what would constitute ‘extremely severe’. The definitions of the items were developed from dictionary definitions to elicit ratings from colloquial understandings rather than injecting expert definitions and biases into the evaluations. After a rater evaluated the first scenario, they were then presented with “a re-written version of the above scenario” (either $S$ or $S^{\prime}$, whichever they did not see first) and performed the same evaluation on this second scenario. To mitigate the potential for ordering effects we randomly counterbalance the order of presentation: three evaluators saw each scenario pair in the order $S$, $S^{\prime}$, and three people saw the scenario in the order $S^{\prime}$, $S$. We later performed a paired sample t-test to assess if there were any ordering biases and found none.
@@ 4 Analysis and Results
For each impact theme below, we will examine (a) which scenarios GPT-4 was adequately able to generate when specifically prompted, (b) the human user study raw scores to understand how these impacts were perceived, and (c) the delta between $S$ and $S^{\prime}$ to evaluate whether humans perceived the policy to be effective at mitigating the designated harm. We begin by examining overall significant trends, then proceed alphabetically by impact theme. Significance refers to results from paired sample t-tests, and we report significance at levels of $p<0.05,0.01$, and $0.001$. All references to a taxonomy of impacts of generative AI in the media environment uncovered by human authors refer to the Kieslich taxonomy (Kieslich, Diakopoulos, and Helberger 2024).
@@
Evaluators believed these issues affected a large portion of society ($M_{S}=4.06;SD_{S}=0.97$) with the second highest average scores for magnitude. They believe the mitigated impacts affected slightly fewer people ($M_{\Delta}=-0.17;SD_{\Delta}=1.01$), and made the impacts less specific to vulnerable populations ($M_{\Delta}=-0.44;SD_{\Delta}=0.83$).
@@
Political impacts across the board tended to be perceived as the most severe of all impact themes ($M_{S}=4.08;SD_{S}=0.76$), with fake news/misinformation having the highest severity ratings of all impact types with a score of $M_{S}=4.56$ ($SD_{S}=0.50$), and manipulation not trailing too far behind ($M_{S}=4.22;SD_{S}=0.92$). On average, transparency legislation seemed to mitigate the severity ($M_{\Delta}=-0.56$***; $SD_{\Delta}=0.86;p<0.001$) of political impacts and even more so in fake news/misinformation ($M_{\Delta}=-0.78$*; $SD_{\Delta}=0.63;p<0.05$) and manipulation ($M_{\Delta}=-0.67$*; $SD_{\Delta}=0.82;p<0.05$). One scenario portraying fake news detailed a fabricated political scandal concocted by an AI news generator and widely disseminated to the public that ultimately led the political candidate to drop out of the race. In the transparency mitigated scenario, an “AI-generated” tag was attached to the piece, and news consumers all paid attention to the human-verified fact checking of the inaccuracies of the purported scandal.
Scenarios detailing political impacts were seen as very plausible ($M_{S}=4.11;SD_{S}=0.97$), second only to labor. They also notably were one of the only impact themes for which the transparency legislation had a perceived significant impact on the specificity to vulnerable populations ($M_{\Delta}=-0.25$*; $SD_{\Delta}=0.68;p<0.05$), indicating that the perceived effect of this legislation may be even more beneficial to those more disparately affected by AI harms.
@@
We were able to generate adequately relevant scenarios for each of the impact types within trustworthiness: discernment of fact and fiction, information chaos, media fatigue, mistrust, and overreliance on AI. These scenarios tended to focus on the relationship between readers and journalists or readers and the news environment at large. They also feature a trajectory towards a lack of trust between society and the news we consume.
@@
5 Discussion and Future Work